Find the cost per ounce of a sunscreen made from 70 oz of lotion that cost $3.44 per ounce and 50 oz of lotion that cost $10.88 per ounce.
A hospital staff mixed a 75% disinfectant solution with a 25% disinfectant solution. How many liters of each were used to make 25 L of a 45% disinfectant solution?
A butcher has some hamburger that is 22% fat and some that is 13% fat. How many pounds of each should be mixed to make 90 lb of hamburger that is 16% fat?
A ferry leaves a harbor and travels to a resort island at an average speed of 18 mph. On the return trip, the ferry travels at an average speed of 12 mph because of fog. The total time for the trip is 5 h. How far is the island from the harbor?

Answer by stanbon(75887)

Find the cost per ounce of a sunscreen made from 70 oz of lotion that cost $3.44 per ounce and 50 oz of lotion that cost $10.88 per ounce.
---
Value 70*3.44+50*10.88 = 784.80 for 120 ounces
---
Cost per ounce = Value/amount = 784.80/120 = $6.54 per ounce.
----------------------------------------------------------------------
A hospital staff mixed a 75% disinfectant solution with a 25% disinfectant solution. How many liters of each were used to make 25 L of a 45% disinfectant solution?
----
active + active = active
0.75x + 0.25(25-x) = 0.45*25
Multiply thru by 100 to get:
75x + 25*25-25x = 45*25
50x = 20*25
x = 10 L (amt of 75% solution needed in the mixture)
---
25-x = 15 L (amt of 25% solution needed in the mixture)
==========================================================
A butcher has some hamburger that is 22% fat and some that is 13% fat. How many pounds of each should be mixed to make 90 lb of hamburger that is 16% fat?
---
fat + fat = fat
0.22x + 0.13(90-x) = 0.16*90
Note: Same procedure as in the previous problem.
================================================

A ferry leaves a harbor and travels to a resort island at an average speed of 18 mph. On the return trip, the ferry travels at an average speed of 12 mph because of fog. The total time for the trip is 5 h. How far is the island from the harbor?
---
h to island DATA:
rate = 18 mph ; distance = x miles ; time = d/r = x/18 hrs.
---------------------
island to h DATA:
rate = 12 mph ; distance = x miles ; time = d/r = x/12 hrs
------
Equation:
time + time = 5 hrs.
x/18 + x/12 = 5
Multiply thru by 36:
2x + 3x = 5*36
5x = 5*36
x = 36 miles
===================
Cheers,
Stan H.
=================